# A Necessary Step toward Faithfulness: Measuring and Improving Consistency in Free-Text Explanations

### v1

@@ Abstract
Faithful free-text explanations are important to ensure transparency in high-stakes AI decision-making contexts, but they are challenging to generate by language models and assess by humans. In this paper, we present a measure for Prediction-EXplanation (PEX) consistency, by extending the concept of weight of evidence. This measure quantifies how much a free-text explanation supports or opposes a prediction, serving as an important aspect of explanation faithfulness. Our analysis reveals that more than 62% explanations generated by large language models lack this consistency. We show that applying direct preference optimization improves the consistency of generated explanations across three model families, with improvement ranging from 43.1% to 292.3%. Furthermore, we demonstrate that optimizing this consistency measure can improve explanation faithfulness by up to 9.7%.
@@ 1 Introduction
Explainable AI systems are those that can specify the relationship between an output or prediction and the deductive or nomological process that led to that prediction through logically consistent and empirically grounded means Woodward and Ross (2003).  [^1] Other forms of explanations exist, such as those that aim to provide useful information for better decision making Kayser et al. (2024); Han et al. (2023) or debugging Ribeiro et al. (2016); Lundberg and Lee (2017). Such explanations are often called “faithful” in the sense that they reflect a model’s true reasoning process Jacovi and Goldberg (2020); Lyu et al. (2024), and faithfulness is often seen as crucial for trustworthiness, transparency, and accountability.
@@
Figure 1: Example of GPT-4 model generating explanations for truthful or deceptive prediction about a hotel review’s authenticity. While each explanation appears plausible on its own, they are inconsistent: the explanation “use a lot of positive language” is used to support both truthful and deceptive predictions.
In this paper, we aim to improve model faithfulness by optimizing the consistency of an explanation with its explanandum (e.g., prediction). If an explanation is to be a faithful presentation of a deductive of nomological process, then it must hold that an explanation $e$ cannot simultaneously support a prediction $y$ and its negation $\lnot y$. For example, in Figure 1, we see that a model uses virtually the same explanation (“use a lot of positive language’) to argue in favor of a hotel review being both authentic and deceptive in this opinion spam classification task.
Our work proceeds in three steps. First, we ask how consistent large language model (LLM)-generated explanations are in practice. To do this, we introduce a measure of consistency, PEX (for Prediction-EXplanation consistency), which can be applied to models that provide probabilities over their generated outputs. PEX leverages the weight of evidence framework Melis et al. (2021); Wod (1985) to quantify how much an explanation speaks in favor of (vs against) a prediction (§ 3.1, § 3.2). On two datasets—TripAdvisor hotel review and Amazon product review spam detection Ott et al. (2013); Hussain et al. (2020)—We find that 62%–86% of explanations from Llama-2 Touvron et al. (2023), Mistral Mistral (2023), and Yi-1.5 Young et al. (2024) are inconsistent, logically implying that they are also unfaithful (§ 6.1).
@@ 2 Related Work
We do not assume access to ground-truth explanations for evaluation, reflecting real-world scenarios; therefore white-box faithfulness evaluation Zhou et al. (2022b); Chen et al. (2018) is not applicable to our problem. While robustness Alvarez-Melis and Jaakkola (2018) and perturbation-based evaluation methods Samek et al. (2016) have been applied in the vision domain, their applicability to NLP is limited due to the discrete nature of language inputs, and removing a single word can render a sentence meaningless Lyu et al. (2024). Simulatability-based methods have been employed to evaluate explanation Hase and Bansal (2020); Doshi-Velez and Kim (2017). This method is recommended to evaluate faithfulness Lyu et al. (2024), as the more faithful an explanation is, the more information it should contain about the model’s decision mechanism, and thus the easier it would be for an external simulator, to predict the model’s behavior based on the explanation. The assumption is that if an explanation leads to a different prediction than that made by the model it explains, then it is unfaithful Jacovi and Goldberg (2020). Pruthi et al. (2022) propose a framework for evaluating explanations by measuring accuracy gains in a student model trained to mimic a teacher. This framework is designed for system-level evaluation rather than individual-level, as the student model requires more than one examples to learn from.
@@ 3.1 Problem: Inconsistent Explanations
Our first goal is to develop a measure of whether a model’s explanations are consistent in justifying its predictions $\bm{a}$ over an its negation $\lnot\bm{a}$, as a faithful explanation cannot simultaneously support both predictions Miller (2019); Lipton (1990). We observe that language model generated explanations can be inconsistent, e.g. in Figure 1, the explanation “use a lot of positive language” supports both the truthful and deceptive hypotheses, failing to distinguish why the model predicted answer $\bm{a}$ instead of the alternative prediction $\lnot\bm{a}$. As a result, the explanation is not consistent and, therefore, not faithful. We focus on generating explanations to justify opinion spam detection Ott et al. (2011); Hussain et al. (2020). We use this task because it requires minimal external knowledge while allowing for the generation of meaningful explanations.
We focus on language models that can answer questions, where a model $M$ takes a question $\bm{q}$ as input and generates an answer $\bm{a}$ as output with probability $M(\bm{a}\mid\bm{q})$. Additionally, we assume that upon request (e.g. prompting or other mechanisms) that the same model $M$ can generate a potentially faithful explanation of its prediction. We denote this as $M(\bm{e}\mid Q(\bm{q},\bm{a}))$, where $Q$ is an appropriate prompt, and explanation $\bm{e}=(e_{1},e_{2},...,e_{n})$ is a sequence of words. We use one-shot example to prompt the model to generate explanation for its prediction. The explanation usually contains 2-3 rationales. The explanations in the one-shot example are generated by GPT-4 (detailed in § A.1).
@@ 3.2 Measuring Prediction-Explanation (PEX) Consistency
We define explanation inconsistency as failing to explicate why the model predicted $\bm{a}$ instead of its negation $\lnot\bm{a}$. To quantify this, we introduce Prediction-EXplanation consistency (PEX), which measures how well an explanation supports a given prediction compared to its negation. We build on the weight of evidence framework Melis et al. (2021); Wod (1985), extending it to to compute probabilities when evidence is given as sequences of words. The weight of evidence measures the extent to which an explanation supports or undermines a hypothesis, making it a reasonable measure of explanation consistency.
@@
To measure the PEX consistency of explanation $\bm{e}$ for a given question $\bm{q}$ and model prediction $\bm{a}$, we compute a score to compare the likelihood of model $M$ generating $\bm{e}$ under different predictions:
@@
where $\lnot\bm{a}$ is a negation from $\bm{a}$, the text prompt $Q$ and explanation generation $M(\bm{e}\mid Q(\bm{q},\bm{a}))$ are formatted in § 3.1. We estimate the conditional probability $M(\bm{e}\mid Q(\bm{q},\bm{a}))$ using the chain rule:
@@
If we view $M(\bm{e}\mid Q(\bm{q},\bm{a}))$ as $M(\bm{e}\mid\bm{q},\bm{a})$ by dropping the text format $Q$, PEX consistency $C(\bm{e})$ can also be computed using Baye’s rule as:
@@
where $\log\frac{M(\bm{a}\mid\bm{q},\bm{e})}{M(\lnot\bm{a}\mid\bm{q},\bm{e})}$ is the posterior log-odds ratio of between prediction $\bm{a}$ and $\lnot\bm{a}$ conditioned on question $\bm{q}$ and explanation $\bm{e}$. $\log\frac{M(\bm{a}\mid\bm{q})}{M(\lnot\bm{a}\mid\bm{q})}$ is the prior log-odds ratio, where $M(\bm{a}\mid\bm{q})$ is defined in § 3.1. To compute $M(\bm{a}\mid\bm{q},\bm{e})$ using a language model $M$, we format the prompt $Q^{\prime}$ as: $Q^{\prime}(\bm{q},\bm{e})$ = Is this review truthful or deceptive? Review: {review}. Analysis: $\bm{e}$. Then, we compute adjusted PEX consistency $C^{\prime}(\bm{e})$ as:
@@
We assume access to language models that provide output probabilities e.g. $M(\bm{a}\mid Q^{\prime}(\bm{q},\bm{e}))$, to compute PEX consistency (Eq 4). Pretrained language models, including LLama-2 and Mistra, do not perform well on opinion spam detection, with prediction accuracy close to random guessing. Thus we fine-tune the models on the training dataset using supervised fine-tuning training (SFT) to have better estimation of $M(\bm{a}\mid\bm{q})$. This is achieved by using the maximum-likelihood estimation (MLE) objective:
@@ 4 Explanation Faithfulness Evaluation
To evaluate the efficacy of optimizing PEX consistency (§ 3) in generating more faithful explanations, we construct a proxy measure of faithfulness by assuming that a more faithful explanation should be more useful for an external observer to predict a model’s behavior (Lyu et al., 2024, see Section 3.4.2). Using this idea, we adopt a simulatability-based explanation faithfulness evaluation method Pruthi et al. (2022), as shown in Figure 3. In this framework, a teacher model generates a prediction and explanation, and a student model is trained to simulate the teacher’s model prediction. Explanations are available to the student model during training, but are not available during testing to avoid label leakage. The student model is models instead of humans, as human experiments are costly and humans’s preconceived notions may impair their ability to simulate the models accurately.
@@
where $R_{n}$ is the $n$-th word of the text response $R(\bm{a},\bm{e})$, and $R_{<n}$ is the first $n-1$ words of the response. Dataset $D_{T}$ consisting of samples in the form of $(\bm{q},\bm{a}_{T},\bm{e}_{T})$, where $\bm{a}_{T}=M_{T}(\bm{q})$ is the answer predicted by the teacher model $M_{T}$, and explanation $\bm{e}_{T}$ is generated by $M_{T}$ to justify its prediction $\bm{a}_{T}$ to the question $\bm{q}$.
@@ 5 Experimental Setup
We generate explanations using the fine-tuned models on the validation split to compute PEX consistency score statistics (§ 3.2). To evaluate explanation faithfulness, we also use the generated explanations on the validation split to serve as teacher model to fine-tune student model (§ 4). We report student model performance on the test split, evaluated against teacher predictions.
@@
We train student models  [^2] Our preliminary experiments do not suggest that using training-free in-context learning is fruitful. using $k$ random examples from the validation split, where each example includes both the teacher model’s prediction and its generated explanation (§ 4). We perform a five-pass training for $k=10$ and $k=20$. We use small values of $k$ because the student model can learn effectively from the prediction label alone when $k$ is large. The student model shares the same architecture as the teacher model but is not fine-tuned on the training split. Further details are provided in § A.2.
@@ 6.1 How consistent are the explanations generated by large language models?
Figure 4 shows the distribution of PEX consistency scores (§ 3.2) for explanations generated by the Mistral, Llama-2 and Yi-1.5 models on two opinion spam datasets. According to Wod (1985); Melis et al. (2021), weight of evidence scores above 2 are considered consistent. For the Mistral model, we see that 85.6% of explanations are inconsistent (with PEX score $<$ 0), 75.4% for the Llama-2 model, and 61.7% for the Yi-1.5 model. Since inconsistent explanations do not support the model’s predictions effectively, they pose a challenge for humans trying to understand how the model arrives at its decisions based on those explanations.
Figure 4: Prediction-EXplanation consistency (PEX) score distribution for different large language models. Explanations with PEX score higher than 2 are considered consistent Wod (1985). Takeaway: models can generate 62%-86% inconsistent explanations.
6.2 Can the consistency of LLM-generated explanations be improved?
(a) TripAdvisor dataset
(b) Amazon dataset
We use direct preference optimization (DPO) to enhance explanation consistency (§ 3.4), the results are illustrated in Figure 5. For the Mistral model, DPO improves consistency by 2.8 to 5.1 points compared to explanations generated from the supervised fine-tuning (SFT) model (§ 3.2). DPO also improves 2.6 to 6.6 points for the Llama-2 model, and 1.8 to 3.8 points for the Yi-1.5 model. These results demonstrate the effectiveness of DPO in improving the PEX consistency of the generated explanations, by learning the patterns that contribute to more consistent explanations. For the Mistral model on the TripAdvisor dataset, the average consistency scores of DPO-generated explanations remain below zero. This arises from the upper bound imposed by the sampled explanations being used to construct preference data: the highest consistency score among DPO-sampled explanations for this model averages -2.1 across the reviews, which is lower than the Llama-2 model’s highest score of 4.4. Similarly, the highest consistency score among DPO-sampled explanations for Mistral model averages -2.8 on the Amazon dataset, which is lower than the Yi-1.5 model’s highest score of 0.3.
6.3 Are explanations optimized for consistency also more faithful?
@@
Figure 6: Direct preference optimization (DPO) generates better explanations: The Mistral SFT (supervised fine-tuning) model predicts the TripAdvisor review as truthful; however, its explanation includes the phrase “mentions an alternative hotel,” which is often associated with deceptive prediction, leading to inconsistency. In contrast, the DPO model produces more faithful explanations with a higher consistency score.

### v2

@@ Abstract
Faithful free-text explanations are important to ensure transparency in high-stakes AI decision-making contexts, but they are challenging to generate by language models and assess by humans. In this paper, we present a measure for Prediction-EXplanation (PEX) consistency, by extending the concept of weight of evidence. This measure quantifies how much a free-text explanation supports or opposes a prediction, serving as an important aspect of explanation faithfulness. Our analysis reveals that more than 62% explanations generated by large language models lack this consistency. We show that applying direct preference optimization improves the consistency of generated explanations across three model families, with improvement ranging from 43.1% to 292.3%. Furthermore, we demonstrate that optimizing this consistency measure can improve explanation faithfulness by up to 9.7%.111 Our code is publicly released at https://github.com/lingjunzhao/PEX_consistency.
@@ 1 Introduction
Explainable AI systems are those that can specify the relationship between an output or prediction and the deductive or nomological process that led to that prediction through logically consistent and empirically grounded means Woodward and Ross (2003).222 Other forms of explanations exist, such as those that aim to provide useful information for better decision making Kayser et al. (2024); Han et al. (2023) or debugging Ribeiro et al. (2016); Lundberg and Lee (2017). Such explanations are often called “faithful” in the sense that they accurately reflect a model’s true reasoning process Jacovi and Goldberg (2020); Lyu et al. (2024), and faithfulness is often seen as crucial for trustworthiness, transparency, and accountability.
@@
Figure 1: Example of GPT-4 model generating explanations for truthful or deceptive prediction about a hotel review’s authenticity. While each explanation appears plausible on its own, they are inconsistent: both the truthful and deceptive explanations rely on the same evidence “use a lot of positive language”.
In this paper, we aim to improve model faithfulness by optimizing the consistency between an explanation and its explanandum (e.g., a model prediction). Following Miller (2019)’s criterion, a necessary condition for faithful explanations is that an explanation should contrastively refute the negation of the explanandum. Therefore, for an explanation to faithfully represent a deductive or nomological process, it cannot simultaneously support a prediction $y$ and its negation $\lnot y$. We formalize this notion as inconsistency. For example, in Figure 1, a model uses essentially the same explanation (“use a lot of positive language’) to argue in favor of a hotel review being both authentic and deceptive in this opinion spam classification task.
Our work proceeds in three steps. First, we ask how consistent large language model (LLM)-generated explanations are in practice. To do this, we introduce a measure of consistency, PEX (for Prediction-EXplanation consistency), which can be applied to models that provide probabilities over their generated outputs. PEX leverages the weight of evidence framework Melis et al. (2021); Good (1985) to quantify how much an explanation speaks in favor of (vs against) a prediction (§ 3.1, § 3.2). On two datasets—TripAdvisor hotel review and Amazon product review spam detection Ott et al. (2013); Hussain et al. (2020)—we find that 62%–86% of explanations from Llama-2 Touvron et al. (2023), Mistral Mistral (2023), and Yi-1.5 Young et al. (2024) are inconsistent, logically implying that they are also unfaithful (§ 6.1).
@@ 2 Related Work
We do not assume access to ground-truth explanations for evaluation, reflecting real-world scenarios; therefore white-box faithfulness evaluation Zhou et al. (2022b); Chen et al. (2018) is not applicable to our problem. While robustness Alvarez-Melis and Jaakkola (2018) and perturbation-based evaluation methods Samek et al. (2016) have been applied in the vision domain, their applicability to NLP is limited due to the discrete nature of language inputs, and removing a single word can render a sentence meaningless Lyu et al. (2024). Simulatability-based methods have been employed to evaluate explanation Hase and Bansal (2020); Hase et al. (2020); Doshi-Velez and Kim (2017). This method is recommended to evaluate faithfulness Lyu et al. (2024), as the more faithful an explanation is, the more information it should contain about the model’s decision mechanism, and thus the easier it would be for an external simulator, to predict the model’s behavior based on the explanation. The assumption is that if an explanation leads to a different prediction than that made by the model it explains, then it is unfaithful Jacovi and Goldberg (2020). Pruthi et al. (2022) propose a framework for evaluating explanations by measuring accuracy gains in a student model trained to mimic a teacher. This framework is designed for system-level evaluation rather than individual-level, as the student model requires more than one examples to learn from.
@@ 3.1 Problem: Inconsistent Explanations
Our first goal is to develop a measure of whether a model’s explanations are consistent in justifying its predictions $\bm{a}$ over an its negation $\lnot\bm{a}$, as a faithful explanation cannot simultaneously support both predictions Lipton (1990); Miller (2019); Brassard et al. (2024). We observe that language model generated explanations can be inconsistent, e.g. in Figure 1, the explanation “use a lot of positive language” supports both the truthful and deceptive hypotheses, failing to distinguish why the model predicted answer $\bm{a}$ instead of the alternative prediction $\lnot\bm{a}$. As a result, the explanation is not consistent and, therefore, not faithful. We focus on generating explanations to justify opinion spam detection Ott et al. (2011); Hussain et al. (2020). We use this task because it requires minimal external knowledge while allowing for the generation of meaningful explanations.
We focus on language models that can answer questions, where a model $M$ takes a question $\bm{q}$ as input and generates an answer $\bm{a}$ as output with probability $M(\bm{a}\mid\bm{q})$. Additionally, we assume that upon request (e.g. prompting or other mechanisms) that the same model $M$ can generate a potentially faithful explanation of its prediction. We denote this as $M(\bm{e}\mid Q(\bm{q},\bm{a}))$, where $Q$ is an appropriate prompt, and explanation $\bm{e}=(e_{1},e_{2},...,e_{n})$ is a sequence of words. We use a one-shot example to prompt the model to generate explanation for its prediction. The explanation usually contains 2-3 rationales. The explanations in the one-shot example are generated by GPT-4 (detailed in § A.1).
@@ 3.2 Measuring Prediction-Explanation (PEX) Consistency
We define explanation inconsistency as failing to explicate why the model predicted $\bm{a}$ instead of its negation $\lnot\bm{a}$. To quantify this, we introduce Prediction-EXplanation consistency (PEX), which measures how well an explanation supports a given prediction compared to its negation. We build on the weight of evidence framework Melis et al. (2021); Good (1985), extending it to to compute probabilities when evidence is given as sequences of words. The weight of evidence measures the extent to which an explanation supports or undermines a hypothesis, making it a reasonable measure of explanation consistency.
@@
To measure the PEX consistency of explanation $\bm{e}$ for a given question $\bm{q}$ and model prediction $\bm{a}$, we compute a score to compare the likelihood of model $M$ generating $\bm{e}$ under different predictions, thereby quantifying how well the explanation supports the given prediction $\bm{a}$ over its negation $\lnot\bm{a}$:
@@
where the text prompt $Q$ and explanation generation $M(\bm{e}\mid Q(\bm{q},\bm{a}))$ are formatted in § 3.1. We estimate the conditional probability $M(\bm{e}\mid Q(\bm{q},\bm{a}))$ using the chain rule:
@@
If we view $M(\bm{e}\mid Q(\bm{q},\bm{a}))$ as $M(\bm{e}\mid\bm{q},\bm{a})$ by dropping the text format $Q$, PEX consistency $C(\bm{e})$ can also be computed using Bayes’ rule as:
@@
where $\log\frac{M(\bm{a}\mid\bm{q},\bm{e})}{M(\lnot\bm{a}\mid\bm{q},\bm{e})}$ is the posterior log-odds ratio between prediction $\bm{a}$ and $\lnot\bm{a}$ conditioned on question $\bm{q}$ and explanation $\bm{e}$. $\log\frac{M(\bm{a}\mid\bm{q})}{M(\lnot\bm{a}\mid\bm{q})}$ is the prior log-odds ratio, where $M(\bm{a}\mid\bm{q})$ is defined in § 3.1. To compute $M(\bm{a}\mid\bm{q},\bm{e})$ using a language model $M$, we format the prompt $Q^{\prime}$ as: $Q^{\prime}(\bm{q},\bm{e})$ = Is this review truthful or deceptive? Review: {review}. Analysis: $\bm{e}$. Then, we compute adjusted PEX consistency $C^{\prime}(\bm{e})$ as:
@@
We assume access to language models that provide output probabilities e.g. $M(\bm{a}\mid Q^{\prime}(\bm{q},\bm{e}))$, to compute PEX consistency (Eq 4). Pretrained language models, including LLama-2 and Mistral, do not perform well on opinion spam detection, with prediction accuracy close to random guessing. Thus we fine-tune the models on the training dataset using supervised fine-tuning training (SFT) to have better estimation of $M(\bm{a}\mid\bm{q})$. This is achieved by using the maximum-likelihood estimation (MLE) objective:
@@ 4 Explanation Faithfulness Evaluation
To evaluate the efficacy of optimizing PEX consistency (§ 3) in generating more faithful explanations, we construct a proxy measure of faithfulness by assuming that a more faithful explanation should be more useful for an external observer to predict a model’s behavior (Lyu et al., 2024, see Section 3.4.2). Using this idea, we adopt a simulatability-based explanation faithfulness evaluation method Pruthi et al. (2022), as shown in Figure 3. In this framework, a teacher model generates a prediction and explanation, and a student model is trained to simulate the teacher’s model prediction. Explanations are available to the student model during training, but are not available during testing to avoid label leakage. We use language models as student models for our evaluation. While human evaluation is valuable, it is prohibitively expensive at scale, and may be confounded by humans’ preconceived notions—especially when conflating explanation faithfulness with plausibility. While language models also reflect biases, they offer a consistent and scalable means of evaluation.
@@
where $R_{n}$ is the $n$-th word of the text response $R(\bm{a},\bm{e})$, and $R_{<n}$ is the first n1n-1 words of the response. Dataset $D_{T}$ consisting of samples in the form of $(\bm{q},\bm{a}_{T},\bm{e}_{T})$, where $\bm{a}_{T}=M_{T}(\bm{q})$ is the answer predicted by the teacher model $M_{T}$, and explanation $\bm{e}_{T}$ is generated by $M_{T}$ to justify its prediction $\bm{a}_{T}$ to the question $\bm{q}$. We generate the answer before the explanation for the student model, rather than after, because our preliminary experiments showed similar performance for both strategies. However, generating explanations at test time is computationally expensive. The answer-first strategy mitigates this overhead by enabling evaluation based on the likelihood of truthful or deceptive predictions without requiring explanation generation during inference.
@@ 5 Experimental Setup
We generate explanations using the fine-tuned models on the validation split to compute PEX consistency score statistics (§ 3.2). To evaluate explanation faithfulness, we also use the generated explanations on the validation split to serve as the teacher model to fine-tune a student model (§ 4). We report student model performance on the test split, evaluated against teacher predictions.
@@
We train student models333 Our preliminary experiments do not suggest that using training-free in-context learning is fruitful. using $k$ random examples from the validation split, where each example includes both the teacher model’s prediction and its generated explanation (§ 4). We perform a five-pass training for $k=10$ and $k=20$. We use small values of $k$ because the student model can learn effectively from the prediction label alone when $k$ is large. The student model shares the same architecture as the teacher model but is not fine-tuned on the training split. Further details are provided in § A.3.
@@ 6.1 How consistent are the explanations generated by large language models?
Consistency distribution.
Figure 4 shows the distribution of PEX consistency scores (§ 3.2) for explanations generated by the Mistral, Llama-2 and Yi-1.5 models on two opinion spam datasets. According to Good (1985); Melis et al. (2021), weight of evidence scores above 2 are considered consistent. For the Mistral model, we see that 85.6% of explanations are inconsistent (with PEX score << 0), 75.4% for the Llama-2 model, and 61.7% for the Yi-1.5 model. Since inconsistent explanations do not support the model’s predictions effectively, they pose a challenge for humans trying to understand how the model arrives at its decisions based on those explanations.
Figure 4: Prediction-EXplanation consistency (PEX) score distribution for different large language models. Explanations with PEX score higher than 2 are considered consistent Good (1985). Takeaway: models can generate 62%-86% inconsistent explanations.
Sensitivity to prompt format.
To assess whether the PEX consistency score is sensitive to different prompt formats, we use 3 different prompt formats (detailed in § A.2) to compute PEX consistency score and compute Kendall rank correlations between PEX consistency scores obtained from each pair of prompt versions. We find that PEX scores are relatively robust to variations in prompt format, as shown in Table 1.
Model V1 vs V2 V1 vs V3 V2 vs V3
Llama-2 (0.42–0.52) (0.33–0.44) (0.52–0.60)
Mistral (0.21–0.33) (0.48–0.56) (0.38–0.48)
Yi-1.5 (0.37–0.47) (0.50–0.58) (0.40–0.50)
Table 1: Kendall rank correlations of PEX consistency scores across prompt versions on the validation set. Ranges indicate 90% bootstrap confidence intervals. Correlations are moderate (: mean $\geq$0.26) to strong (: mean $\geq$0.49). Takeaway: PEX scores are relatively robust to variations in prompt format.
@@
6.2 Can the consistency of LLM-generated explanations be improved?
We use direct preference optimization (DPO) to enhance explanation consistency (§ 3.4), the results are illustrated in Figure 5. For the Mistral model, DPO improves consistency by 2.8 to 5.1 points compared to explanations generated from the supervised fine-tuning (SFT) model (§ 3.2). DPO also improves 2.6 to 6.6 points for the Llama-2 model, and 1.8 to 3.8 points for the Yi-1.5 model. These results demonstrate the effectiveness of DPO in improving the PEX consistency of the generated explanations, by learning the patterns that contribute to more consistent explanations.
For the Mistral model on the TripAdvisor dataset, the average consistency scores of DPO-generated explanations remain below zero. This arises from the upper bound imposed by the sampled explanations being used to construct preference data: the highest consistency score among DPO-sampled explanations for this model averages -2.1 across the reviews, which is lower than the Llama-2 model’s highest score of 4.4. Similarly, the highest consistency score among DPO-sampled explanations for Mistral model averages -2.8 on the Amazon dataset, which is lower than the Yi-1.5 model’s highest score of 0.3.
(a) TripAdvisor dataset
(b) Amazon dataset
6.3 Are explanations optimized for consistency also more faithful?
Figure 6: Direct preference optimization (DPO) generates better explanations: The Mistral SFT (supervised fine-tuning) model predicts the TripAdvisor review as truthful; however, its explanation includes the phrase mentions an alternative hotel, which is often associated with deceptive prediction (e.g. Comparison to other hotels: The reviewer mentions other hotels that are "better" or have "better service"), leading to inconsistency. In contrast, the DPO model produces more faithful explanations with a higher consistency score.